FIND THE MAXIMUM AND THE MINIMUM VALUES OF
3X^4 - 2X^3 - 6X^2 + 6X + 1 IN THE INTERVAL [0,2] IF THEY EXIST
1
Expert's answer
2020-02-19T05:12:02-0500

FIND THE MAXIMUM AND THE MINIMUM VALUES OF

3 X 4- 2 X 3 - 6 X 2 + 6 X + 1 IN THE INTERVAL [0,2] IF THEY EXIST

f(x)=3 X 4 - 2 X 3 - 6 X 2+ 6 X + 1

f'(x)=12 x 3-6 x 2-12 x+6

12 x 3-6 x 2-12 x+6=0

6(2 x 3-x 2-2 x+1)=0

by finding synthetic division

the roots are

x=1,-1,1/2

critical points are 0,1/2,1,2

f(1)=3-2-6+6+1

f(1)=2

f(1/2)=3(1/2) 4-2(1/2)3-6(1/2)2+6(1/2)+1

=39/16

f(0)=1

f(2)=3(2 4)-2(x 3)-6(22)+6(2)+1

=21

minimum is at (0,1),minimum is at (1,2), maximum is at (1/2,39/16), maximum is at (2,21).

Absolute maximum=21.

Absolute minimum=1,

local minima=2,

local maxima=39/16,

the domain is (- infinity ,+ infinity).